OpenStudy (anonymous):

solve by using LCD : 1/10 +1/y=1/5

OpenStudy (lgbasallote):

hint: the LCD is 10 does that help?

OpenStudy (anonymous):

i know it would be 10y but im stuck after that

OpenStudy (lgbasallote):

yeah. 10y sorry... multiply everything by 10y \[(\frac 1{10} \times 10y) + (\frac 1y \times 10y) = (\frac 15 \times 10y)\] do you know what to do now?

OpenStudy (anonymous):

yeah, got it thanks